# LLM Unlearning with LLM Beliefs

### v1

@@ Abstract
Large language models trained on vast corpora inherently risk memorizing sensitive or harmful content, which may later resurface in their outputs. Prevailing unlearning methods generally rely on gradient ascent and its variants to lower the probability of specific target responses. However, we find that this strategy induces a critical side effect: probability mass is redistributed into high-likelihood regions, often corresponding to semantically related rephrasings of the targets. We refer to this as the squeezing effect, which explains why many methods yield merely spurious unlearning, a problem further obscured by automated metrics (e.g., ROUGE, truth ratio) that misreport actual success. To address this, we propose a bootstrapping (BS) framework that explicitly links the squeezing effect with the model’s own high-confidence generations, namely its model beliefs. Since model beliefs inherently capture the very high-likelihood regions where probability mass is squeezed, incorporating them into the unlearning objective directly counters the squeezing effect. By jointly suppressing both target responses and model beliefs, BS-T (token) attenuates high-probability tokens, whereas BS-S (sequence) removes entire high-confidence generations, together achieving more thorough forgetting while preserving utility. Extensive experiments across diverse benchmarks with various model families confirm the effectiveness of our approach.
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable success in generation and comprehension across diverse applications (Hadi et al., 2023), yet their deployment requires careful auditing to prevent leakage of private, illegal, or misleading information. A common practice is the “report then remove” pipeline (Geng et al., 2025), where harmful behaviors are first identified and then eliminated by model owners. Recently, LLM unlearning (Yao et al., 2024a; Maini et al., 2024; Zhang et al., 2024) emerges as a more principled solution, aiming to directly erase harmful parameterizations from the model itself. Compared with alternatives such as harmful content detectors or in-context defenses (Shi et al., 2024; Pawelczyk et al., 2024), unlearning is less vulnerable to circumvention, jailbreaks, or re-training attacks for open-source LLMs (Lynch et al., 2024).
@@
To achieve unlearning, many studies employ gradient ascent (GA) (Eldan & Russinovich, 2023; Yao et al., 2024b), which inverts the conventional gradient descent process by maximizing the negative log-likelihood (NLL) of to-be-unlearned data so as to erase their influence from the parameters. However, directly applying GA can notably degrades overall performance (Wang et al., 2025a; d; Zhu et al., 2025), limiting practical utility. Consequently, subsequent works pursue refinements, either by improving GA itself (e.g., NPO (Zhang et al., 2024) and WGA (Wang et al., 2025b)), or by incorporating regularization (e.g., GradDiff (Maini et al., 2024)) to better preserve utility. For detailed related works about machine unlearning and LLM unlearning, please refer to Appendix B.
@@
Despite recent refinements, GA-based methods display an intuitive yet underexplored failure mode: unlearned models continue to generate semantically rephrased outputs that retain the knowledge intended for removal, leading to only superficial forgetting. This spurious unlearning is evident to humans but poorly captured by widely used metrics such as ROUGE and perplexity (Maini et al., 2024; Li et al., 2024b), which evaluate surface similarity rather than whether harmful knowledge remains encoded. To uncover such cases, we employ LLM-based evaluation as an auxiliary probe, which reveals that models judged successful by classical metrics may still leak targeted knowledge (cf. Section 3.1). Motivated by this evidence, we in Section 3.2 analyze the mechanism behind spurious unlearning: GA lowers the likelihood of the target response, yet softmax normalization redistributes probability mass to other tokens and sequences, concentrating on high-probability neighborhoods that correspond to paraphrases or closely related continuations (Ren & Sutherland, 2025; Razin et al., 2025). Outputs sampled from these regions thus remain semantically tied to the original target. Figure 1 illustrates this squeezing effect, where suppression of the target response inadvertently elevates related alternatives. This observation suggests a remedy: effective unlearning ought to suppress not only target responses but also the model’s own high-confidence generations—its model beliefs, namely the tokens or sequences it would otherwise predict with highest confidence—thereby preventing probability mass from shifting to semantically similar rephrasings and motivating the framework developed next.
@@
Building on the above insight, we propose a bootstrapping (BS) framework (Yarowsky, 1995), where “bootstrapping” reflects the idea of using the model beliefs as auxiliary unlearning signals. This design extends unlearning beyond fixed target responses and directly counteracts the probability regions into which mass would otherwise be squeezed, thereby enabling more thorough forgetting. Concretely, BS is realized in two forms: BS-token (BS-T) mixes the one-hot label of the target response with the model’s own high-probability token predictions to form a soft target, explicitly suppressing those tokens during training; BS-sequence (BS-S) samples entire high-confidence responses from the model and augments them as additional unlearning data, ensuring that complete harmful continuations are removed rather than merely isolated words (cf. Section 4.1). In both cases, model beliefs are directly built into the loss: the objective penalizes not only the original target but also what the model itself would otherwise most confidently predict, preventing probability mass from “escaping” into semantically similar rephrasings. We further provide theoretical analysis showing how such bootstrapping alleviates the squeezing effect under the learning dynamics framework (cf. Section 4.2). Finally, in Section 5, extensive experiments conducted with OpenUnlearning (Dorna et al., 2025) across multiple benchmarks and models confirm the effectiveness of BS-T and BS-S over prior methods.
@@
  • We propose a bootstrapping-based framework that incorporates model beliefs into the unlearning objective. Instantiated at the token level (BS-T) and sequence level (BS-S), it dynamically suppresses both target responses and high-confidence alternatives. We further provide theoretical analysis showing how BS-T reshapes gradient dynamics and mitigates the squeezing effect.
@@ 2 Preliminaries: From Concepts to Practices
To begin with, we introduce relevant concepts and notations, involving the problem definitions of LLM unlearning in Section 2.1, the existing methods in Section 2.2, as well as the existing evaluations in Section 2.3.
2.1 Problem Definitions
@@
LLM Unlearning. LLMs trained on large datasets ${\mathcal{D}}_{\rm t}$ with parameters $\bm{\theta}_{\mathrm{o}}$} inevitably acquire not only broad capabilities but also harmful or undesirable knowledge that may surface in outputs. LLM unlearning aims to reverse the learning process by adjusting parameters post hoc to remove such knowledge. It relies on an unlearning dataset ${\mathcal{D}}_{\rm u}\subseteq{\mathcal{D}}_{\rm t}$ of prompt–response pairs $({\mathbf{x}}_{\rm u},{\mathbf{y}}_{\rm u})$ to be forgotten, together with a complementary retention dataset ${\mathcal{D}}_{\rm r}$ of pairs $({\mathbf{x}}_{\rm r},{\mathbf{y}}_{\rm r})$, either drawn from ${\mathcal{D}}_{\rm t}\setminus{\mathcal{D}}_{\rm u}$ or constructed independently to specify behaviors to retain. The goal is twofold:
  • Unlearning: the unlearned model with parameters $\bm{\theta}_{\mathrm{u}}$} should assign low likelihood to responses in ${\mathcal{D}}_{\rm u}$ and their rephrasings $\tilde{\mathcal{D}}_{\rm u}$;
  • Retention: for inputs outside $\tilde{\mathcal{D}}_{\rm u}$, its output distribution $\pi_{{\bm{\theta}_{\mathrm{u}}}}(\cdot|{\mathbf{x}})$ should remain close to that of the original model, i.e., $\pi_{{\bm{\theta}_{\mathrm{o}}}}(\cdot|{\mathbf{x}})$.
Achieving both unlearning and retention simultaneously is crucial for reliable deployment but remains challenging, since existing methods often compromise one objective for the other (Zhang et al., 2024; Wang et al., 2025d).
@@ 2.3 Existing Evaluations
Metric-based Evaluations. Most prior work relies on classical metrics, often benchmark-specific. Common choices include Probability and Perplexity (Maini et al., 2024), which measure the likelihood of generating target responses; ROUGE (Lin & Och, 2004), which assesses similarity to the ground truth; QA Accuracy (Li et al., 2024b), which measures the model preference for correct responses; and Extraction Strength (Wang et al., 2025a), which quantifies the degree of knowledge parameterization. These metrics can capture both unlearning and retention, but their failure cases remain largely underexplored, with only a few conceptual studies (Wang et al., 2025a).
Detector- and LLM-based Evaluations. Other studies use task-specific detectors or use LLM-as-a-judge (LaaJ) (Zheng et al., 2023; Chiang & Lee, 2023). Detectors include reward models for retention and harmful-content detectors for safety (Lynch et al., 2024), while LaaJ evaluates whether generated responses still reflect familiarity with unlearned data, such as copyrighted content (Wei et al., 2025). Although less common, LLM-based evaluations often yield more accurate judgments than classical metrics and are later used in this work to reveal spurious unlearning.
@@ 3 Rethinking Existing Works: Failure Modes and Mechanisms
Despite advances in algorithms and evaluations, it remains uncertain whether current unlearning results truly reflect reliable forgetting. Prior studies rarely scrutinize the validity of adopted metrics, casting doubts on the reported gains. This section examines the reliability of such evaluations and uncovers the mechanisms behind apparent successes that, de facto, still preserve forgotten knowledge. Section 3.1 presents case studies that reveal inconsistencies between metric-reported success and human judgment. Section 3.2 further analyzes how NPO-based methods inherently redistribute probability mass into semantically related regions, which explains why models often exhibit only superficial unlearning.
@@ 3.1 Case Studies: Identifying Spurious Unlearning under Misleading Metrics
We first present failure cases where metric-reported success diverges from the actual outcomes manifested in model responses. Our experiments use the TOFU benchmark (Maini et al., 2024), which targets removing private content. We consider GA and NPO, which are widely used baselines underpinning many later works (Yang et al., 2025; Fan et al., 2025). We evaluate the 10% forgetting setup with LLaMA-3.2-1B under greedy decoding, which is stricter than sampling and better highlights failure cases. Results are reported under TOFU-suggested metrics, including Probability, ROUGE-L, and Truth Ratio, where smaller values indicate stronger removal [^1] We adopt inverted Truth Ratio, $\max(0,1-R_{\text{truth}})$, so that lower scores indicate more effective unlearning..
@@
  • Similarity. Echoing Case 2, model responses after unlearning should differ notably from the original ones, thereby preventing privacy leakage or exposure to harmful content. This objective aligns with the unlearning goal in Section 2.1, where seeks to eliminate the associated knowledge rather than merely removing the unlearning corpora.
These two perspectives operationalized as LaaJ prompts, with ratings from 0 (failure) to 5 (success) indicating the unlearning strength. See Appendix F.2 for further details of our LaaJ evaluation.
@@ 3.2 Mechanistic Analysis: The Squeezing Effect behind Spurious Unlearning
In Section 3.1, two distinct failure modes of LLM unlearning are identified. Case 1 has been investigated in prior work (Wang et al., 2025b), in which the inverse likelihood derived from GA gradients leads to degenerate outputs. Here, we shift our attention to Case 2, where models still produce rephrased responses that retain the original semantics. This section aims to uncover the mechanism behind such spurious unlearning, a phenomenon largely overlooked in existing studies.
Our Conjecture. We hypothesize that spurious unlearning arises from a redistribution of probability mass enforced by the softmax constraint. Since the conditional probabilities for a given input must sum to one, lowering the likelihood of the target response $\pi_{{\bm{\theta}}}({\mathbf{y}}_{\rm u}|{\mathbf{x}}_{\rm u})$ inevitably increases the likelihood of some alternative candidates, i.e., $\pi_{{\bm{\theta}}}({\mathbf{y}}|{\mathbf{x}}_{\rm u})$ for ${\mathbf{y}}\neq{\mathbf{y}}_{\rm u}$. This increase typically occurs on high-likelihood regions, where generated responses are semantically similar to the original due to the LLM pre-training generalization. Consequently, the model tends to replace exact matches with semantically related rephrasings, a behavior we term the squeezing effect, borrowing terminology from LLM finetuning (Ren & Sutherland, 2025).
Empirical Verification. To examine our conjecture, we conduct two complementary experiments on TOFU under 10% forget setting. First, we use beam search to sample diverse responses from the original LLM and group them by conditional probability into high-, mid-, and low-likelihood regions (top 20%, 20–60%, and 60–100%). Their semantic overlap with the original targets is then evaluated using LaaJ similarity in Section 3.1, and compared with responses generated by retraining (i.e., standard gold model) and by NPO. The results in Figure 2(a) directly quantify semantic preservation across different likelihood bands and unlearning strategies. Second, we track the log-probability dynamics of these groups during GA and NPO training (Figure 2(b) and 2(c)), which reveal how probability mass is redistributed throughout optimization. From these experiments we derive two key observations:
@@
    Semantic correlation concentrates in high-likelihood regions. As shown in Figure 2(a), responses from the high-likelihood region are consistently judged by LaaJ as most semantically related to the original outputs, whereas mid- and low-likelihood regions exhibit lower similarity. Notably, after unlearning, NPO’s generations remain considerably more semantically related than retrain, with similarity scores only slightly below high-likelihood paraphrases and above the mid-likelihood band. This indicates that spurious unlearning is not a corner case (as in Case 2) but a systematic outcome of NPO: it suppresses exact matches yet retains semantically overlapping responses.
@@
    Probability mass is persistently squeezed into these regions. Figure 2(b) and 2(c) show that both GA and NPO initially amplify the likelihood of high-probability responses when suppressing targets, confirming that mass is redistributed into nearby semantic neighborhoods. Although GA’s aggressive updates eventually degrade the model (Wang et al., 2025b) and diminish this effect, NPO maintains the squeezing pattern in a more stable manner. This persistence explains why NPO often yields surface-level forgetting but continues to expose underlying knowledge through paraphrased outputs, aligning with the limited generalization observed in Case 2.

### v2

@@ Abstract
Large language models trained on vast corpora inherently risk memorizing sensitive or harmful content, which may later resurface in their outputs. Prevailing unlearning methods generally rely on gradient ascent and its variants to lower the probability of specific target responses. However, we find that this strategy induces a critical side effect: probability mass is redistributed into high-likelihood regions, often corresponding to semantically related rephrasings of the targets. We refer to this as the squeezing effect, which explains why many methods yield merely spurious unlearning, a problem further obscured by automated metrics (e.g., ROUGE, truth ratio) that misreport actual success. To address this, we propose a bootstrapping (BS) framework that explicitly links the squeezing effect with the model’s own high-confidence generations, namely its model beliefs. Since model beliefs inherently capture the very high-likelihood regions where probability mass is squeezed, incorporating them into the unlearning objective directly counters the squeezing effect. By jointly suppressing both target responses and model beliefs, BS-T (token) attenuates high-probability tokens, whereas BS-S (sequence) removes entire high-confidence generations, together achieving more thorough forgetting while preserving utility. Extensive experiments on diverse benchmarks confirm the effectiveness of our approach, with code merged to OpenUnlearning.
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable success in generation and comprehension across diverse applications (Hadi et al., 2023; Zhang et al., 2026b), yet their deployment requires careful auditing to prevent leakage of private, illegal, or misleading information. A common practice is the “report then remove” pipeline (Geng et al., 2025), where harmful behaviors are first identified and then eliminated by model owners. Recently, LLM unlearning (Yao et al., 2024a; Zhu et al., 2025; Liao et al., 2026) emerges as a more principled solution, aiming to directly erase harmful parameterizations from the model itself. Compared with alternatives such as harmful content detectors or in-context defenses (Shi et al., 2024; Pawelczyk et al., 2024), unlearning is less vulnerable to circumvention, jailbreaks, or re-training attacks for open-source LLMs (Lynch et al., 2024).
@@
To achieve unlearning, many studies employ gradient ascent (GA) (Eldan and Russinovich, 2023; Yao et al., 2024b), which inverts the conventional gradient descent process by maximizing the negative log-likelihood (NLL) of to-be-unlearned data so as to erase their influence from the parameters. However, directly applying GA can notably degrades overall performance (Wang et al., 2025a; d; Zhu et al., 2025), limiting practical utility. Consequently, subsequent works pursue refinements, either by improving GA itself (e.g., NPO (Zhang et al., 2024) and WGA (Wang et al., 2025b)), or by incorporating regularization (e.g., GradDiff (Maini et al., 2024)) to better preserve utility. For detailed related works about machine unlearning and LLM unlearning, please refer to Appx. B.
@@
Despite recent refinements, GA-based methods display an intuitive yet underexplored failure mode: unlearned models continue to generate semantically rephrased outputs that retain the knowledge intended for removal, leading to only superficial forgetting. This spurious unlearning is evident to humans but poorly captured by widely used metrics such as ROUGE and perplexity (Zhu et al., 2026; Li et al., 2024b; Wang et al., 2025e; Chen et al., 2026), which evaluate surface similarity rather than whether harmful knowledge remains encoded. To uncover such cases, we employ LLM-based evaluation as an auxiliary probe, which reveals that models judged successful by classical metrics may still leak targeted knowledge (cf. §3.1). Motivated by this evidence, we in §3.2 analyze the mechanism behind spurious unlearning: GA lowers the likelihood of the target response, yet softmax normalization redistributes probability mass to other tokens and sequences, concentrating on high-probability neighborhoods that correspond to paraphrases or closely related continuations (Ren and Sutherland, 2025; Razin et al., 2025). Outputs sampled from these regions thus remain semantically tied to the original target. Fig. 1 illustrates this squeezing effect, where suppression of the target response inadvertently elevates related alternatives. This observation suggests a remedy: effective unlearning ought to suppress not only target responses but also the model’s own high-confidence generations—its model beliefs, namely the tokens or sequences it would otherwise predict with highest confidence—thereby preventing probability mass from shifting to semantically similar rephrasings.
@@
Building on the above insight, we propose a bootstrapping (BS) framework (Yarowsky, 1995), where “bootstrapping” reflects the idea of using the model beliefs as auxiliary unlearning signals. This design extends unlearning beyond fixed target responses and directly counteracts the probability regions into which mass would otherwise be squeezed, thereby enabling more thorough forgetting. Concretely, BS is realized in two forms: BS-token (BS-T) mixes the one-hot label of the target response with the model’s own high-probability token predictions to form a soft target, explicitly suppressing those tokens during training; BS-sequence (BS-S) samples entire high-confidence responses from the model and augments them as additional unlearning data, ensuring that complete harmful continuations are removed rather than merely isolated words (cf. §4). In both cases, model beliefs are directly built into the loss: the objective penalizes not only the original target but also what the model itself would otherwise most confidently predict, preventing probability mass from “escaping” into semantically similar rephrasings. We further provide theoretical analysis showing how such bootstrapping alleviates the squeezing effect under the learning dynamics framework (cf. §5). Finally, in §6, extensive experiments conducted with OpenUnlearning (Dorna et al., 2025) across multiple benchmarks and models confirm the effectiveness of BS-T and BS-S over prior methods.
@@
  • We propose a bootstrapping-based framework that incorporates model beliefs into the unlearning objective. Instantiated at the token level (BS-T) and sequence level (BS-S), it dynamically suppresses both target responses and high-confidence alternatives. We further provide theoretical analysis showing how BS reshapes gradient dynamics and mitigates the squeezing effect.
@@ 2 Preliminaries: From Concepts to Practices
2.1 Problem Definition
@@
LLM Unlearning. LLMs trained on large datasets ${\mathcal{D}}_{\rm t}$ with parameters $\bm{\theta}_{\mathrm{o}}$} inevitably acquire not only broad capabilities but also harmful or undesirable knowledge that may surface in outputs. LLM unlearning aims to reverse the learning process by adjusting parameters post hoc to remove such knowledge. It relies on an unlearning dataset ${\mathcal{D}}_{\rm u}\subseteq{\mathcal{D}}_{\rm t}$ of prompt–response pairs $({\mathbf{x}}_{\rm u},{\mathbf{y}}_{\rm u})$ to be forgotten, together with a complementary retention dataset ${\mathcal{D}}_{\rm r}$ of pairs $({\mathbf{x}}_{\rm r},{\mathbf{y}}_{\rm r})$, either drawn from ${\mathcal{D}}_{\rm t}\setminus{\mathcal{D}}_{\rm u}$ or constructed independently to specify behaviors to retain. The goal is twofold: 1) Unlearning: the unlearned model with parameters $\bm{\theta}_{\mathrm{u}}$} should assign low likelihood to responses in ${\mathcal{D}}_{\rm u}$ and their rephrasings $\tilde{\mathcal{D}}_{\rm u}$; 2) Retention: for inputs outside $\tilde{\mathcal{D}}_{\rm u}$, its output distribution $\pi_{{\bm{\theta}_{\mathrm{u}}}}(\cdot|{\mathbf{x}})$ should remain close to that of the original model, i.e., $\pi_{{\bm{\theta}_{\mathrm{o}}}}(\cdot|{\mathbf{x}})$. Achieving both unlearning and retention simultaneously is crucial for reliable deployment but remains challenging, since existing methods often compromise one objective for the other (Zhang et al., 2024; Wang et al., 2025d).
@@ 2.3 Existing Evaluations
Metric-based Evaluations. Most prior work relies on classical metrics, often benchmark-specific. Common choices include Probability and Perplexity (Maini et al., 2024), which measure the likelihood of generating target responses; ROUGE (Lin and Och, 2004), which assesses similarity to the ground truth; QA Accuracy (Li et al., 2024b), which measures the model preference for correct responses; and Extraction Strength (Wang et al., 2025a), which quantifies the degree of knowledge parameterization. These metrics can capture both unlearning and retention, but their failure cases remain largely underexplored, with only a few conceptual studies (Wang et al., 2025a).
Detector- and LLM-based Evaluations. Other studies use task-specific detectors or use LLM-as-a-judge (LaaJ) (Zheng et al., 2023; Chiang and Lee, 2023; Zhang et al., 2026a). Detectors include reward models for retention and harmful-content detectors for safety (Lynch et al., 2024), while LaaJ evaluates whether generated responses still reflect familiarity with unlearned data, such as copyrighted content (Wei et al., 2025). Although less common, LLM-based evaluations often yield more accurate judgments than classical metrics and are later used in this work to reveal spurious unlearning.
@@ 3 Rethinking Existing Works: Failure Modes and Mechanisms
Despite advances in algorithms and evaluations, it remains uncertain whether current unlearning results truly reflect reliable forgetting. Prior studies rarely scrutinize the validity of adopted metrics, casting doubts on the reported gains. This section examines the reliability of such evaluations and uncovers the mechanisms behind apparent successes that, de facto, still preserve forgotten knowledge. §3.1 presents case studies that reveal inconsistencies between metric-reported success and human judgment. §3.2 further analyzes how NPO-based methods inherently redistribute probability mass into semantically related regions, which explains why models often exhibit only superficial unlearning.
@@ 3.1 Case Studies: Identifying Spurious Unlearning under Misleading Metrics
We first present failure cases where metric-reported success diverges from the actual outcomes manifested in model responses. Our experiments use the TOFU benchmark (Maini et al., 2024), which targets removing private content. We consider GA and NPO, which are widely used baselines underpinning many later works (Yang et al., 2025; Fan et al., 2025). We evaluate the 10% forgetting setup with Llama 3.2 1B under greedy decoding, which is stricter than sampling and better highlights failure cases. Results are reported under TOFU-suggested metrics, including Probability, ROUGE-L, and Truth Ratio, where smaller values indicate stronger removal [^1] We adopt inverted Truth Ratio, $\max(0,1-R_{\text{truth}})$, so that lower scores indicate more effective unlearning..
@@
  • Similarity. Echoing Case 2, model responses after unlearning should differ notably from the original ones, thereby preventing privacy leakage or exposure to harmful content. This objective aligns with the unlearning goal in §2.1, where seeks to eliminate the associated knowledge rather than merely removing the unlearning corpora.
These two perspectives operationalized as LaaJ prompts, with ratings from 0 (failure) to 5 (success) indicating the unlearning strength. See Appx. F.2 for further details of our LaaJ evaluation.
@@ 3.2 Mechanistic Analysis: The Squeezing Effect behind Spurious Unlearning
In §3.1, two distinct failure modes of LLM unlearning are identified. Case 1 has been investigated in prior work (Wang et al., 2025b), in which the inverse likelihood derived from GA gradients leads to degenerate outputs. Here, we shift our attention to Case 2, where models still produce rephrased responses that retain the original semantics. This section aims to uncover the mechanism behind such spurious unlearning, a phenomenon largely overlooked in existing studies.
Our Conjecture. We hypothesize that spurious unlearning arises from a redistribution of probability mass enforced by the softmax constraint. Since the conditional probabilities for a given input must sum to one, lowering the likelihood of the target response $\pi_{{\bm{\theta}}}({\mathbf{y}}_{\rm u}|{\mathbf{x}}_{\rm u})$ inevitably increases the likelihood of some alternative candidates, i.e., $\pi_{{\bm{\theta}}}({\mathbf{y}}|{\mathbf{x}}_{\rm u})$ for ${\mathbf{y}}\neq{\mathbf{y}}_{\rm u}$. This increase typically occurs on high-likelihood regions, where generated responses are semantically similar to the original due to the LLM pre-training generalization. Consequently, the model tends to replace exact matches with semantically related rephrasings, a behavior we term the squeezing effect, borrowing terminology from LLM finetuning (Ren and Sutherland, 2025).
Empirical Verification. To examine our conjecture, we conduct two complementary experiments on TOFU under 10% forget setting. First, we use beam search to sample diverse responses from the original LLM and group them by conditional probability into high-, mid-, and low-likelihood regions (top 20%, 20–60%, and 60–100%). Their semantic overlap with the original targets is then evaluated using LaaJ similarity in §3.1, and compared with responses generated by retraining (i.e., standard gold model) and by NPO. The results in Fig. 2(a) directly quantify semantic preservation across different likelihood bands and unlearning strategies. Second, we track the log-probability dynamics of these groups during GA and NPO training (Fig. 2(b) and 2(c)), which reveal how probability mass is redistributed throughout optimization. From these experiments we derive two key observations:
@@
    Semantic correlation concentrates in high-likelihood regions. As shown in Fig. 2(a), responses from the high-likelihood region are consistently judged by LaaJ as most semantically related to the original outputs, whereas mid- and low-likelihood regions exhibit lower similarity. Notably, after unlearning, NPO’s generations remain considerably more semantically related than retrain, with similarity scores only slightly below high-likelihood paraphrases and above the mid-likelihood band. This indicates that spurious unlearning is not a corner case (as in Case 2) but a systematic outcome of NPO: it suppresses exact matches yet retains semantically overlapping responses.
@@
    Probability mass is persistently squeezed into these regions. Fig. 2(b) and 2(c) show that both GA and NPO initially amplify the likelihood of high-probability responses when suppressing targets, confirming that mass is redistributed into nearby semantic neighborhoods. Although GA’s aggressive updates eventually degrade the model (Wang et al., 2025b) and diminish this effect, NPO maintains the squeezing pattern in a more stable manner. This persistence explains why NPO often yields surface-level forgetting but continues to expose underlying knowledge through paraphrased outputs, aligning with the limited generalization observed in Case 2.